# Toward Stable and Consistent Evaluation Results: A New Methodology for Base Model Evaluation

### v1

@@ Abstract
Toward Stable and Consistent Evaluation Results: A New Methodology for Base Model Evaluation
Hongzhi Luan [^1] Equal contributions.,  Changxin Tianthanks: Equal contributions.,  Zhaoxin Huan,  Xiaolu Zhang Kunlong Chen,  Zhiqiang Zhang,  Jun Zhouthanks: Corresponding author. {luanhongzhi.lhz, tianchangxin.tcx, zhaoxin.hzx, yueyin.zxl}@antgroup.com {kunlong.ckl, lingyao.zzq, jun.zhoujun}@antgroup.com
@@ 1 Introduction
Recently, large language models (LLMs) have demonstrated remarkable achievements across various domains Minaee et al. (2024). This has led to the development of numerous high-performing LLMs OpenAI (2024); Aaron Grattafiori and Abhinav Jauhri (2024); Team (2025, 2024a). To evaluate the performance of these models on a wide range of tasks, an increasing number of benchmarks have been open-sourced Guo et al. (2023). These benchmarks provide a comprehensive evaluation of the capabilities of LLMs, guiding the training and improving their weakness. Evaluating large language models is a cornerstone in development of LLMs.
@@ 2 Related Work
Due to the above issues, although many technical reports of open-source LLMs disclose the benchmarks and evaluation results used for base model evaluation Aaron Grattafiori and Abhinav Jauhri (2024); Team (2025, 2024a), the lack of unified evaluation protocol and the incomplete transparency of evaluation details make it challenging to reproduce and align the evaluation results, which may cause significant difficulties for researchers.
@@ 3.1 Open-ended Task
Light-instruction prompt. While prompting techniques have proven effective for LLMs Wei et al. (2023a); Sander Schulhoff and et al. (2024), excessive prompting may confuse base models due to their lack of instruction-tuning Wang and Zhou (2024). We propose a lightweight prompt template, which formulates the input using the following format:
@@ 4 Experiments
In this section, we describe systematic experiments to investigate how BOSE enhances evaluation stability during pre-training and ensures consistency between base and instruct models [^1] All of our experiments are based on OpenCompass OpenCompass (2023), and the related code will be available after the reviewing process..
@@ 4.1.2 Models
Pre-trained models [^3] Since the open-source LLMs do not provide publicly accessible intermediate model weights, we take our pre-trained models for experiments.. These models are collected from our pre-training experiments, encompassing 1B and 2B parameter sizes with different training steps, 96 checkpoints in total. To assess the capability consistency between base models and their instruction-tuned counterparts, we further fine-tune 24 checkpoints using identical instruction data to derive the instruct models.
Open-source base models. These models mainly cover three remarkable LLM families: (1) Llama family Aaron Grattafiori and Abhinav Jauhri (2024): mainly including Llama-3.1-8B/70B, with an extended analysis incorporating Llama-3-8B/70B and Llama-3.2-1B/3B for consistency investigation; (2) Gemma family Team (2024a): Gemma-2-9B/27B; (3) Qwen family Team (2025): Qwen2.5-7B/72B.
@@ 4.1.3 Metric
Consistency with instruct models: measured by the Kendall’s rank correlation between base models’ evaluation scores and corresponding instruct models’ scores.
@@ 4.4 Ablation Study
To further investigate whether all components of the proposed ICLiP methodology are indispensable, we conduct ablation experiments here. The following variant prompt templates are considered as comparative methods: (1) instruct_0shot: instruct prompt with 0-shot, also referred to as the original method, (2) instruct_fewshot: instruct prompt with few-shot, (3) light-instruction_0shot: light-instruction prompt with 0-shot, (4) ICLiP, our proposed method. Here, light-instruction refers to the prompt template used in our proposed method, while instruct prompt is the prompt template typically used for instruct model evaluation. Please refer to Appendix A for further implementation details.
@@ 6 Limitations
Our work aims to address some critical issues in base model evaluation, such as instability during pre-training and lack of consistency with instruct models. There may be some subjective biases in our comparative methods, such as the design of instruction prompt templates. As for the considered models, we use our pre-trained checkpoints and 3 open-source LLM families, and there are more open-source LLMs worth exploring, such as DeepSeek DeepSeek-AI (2024), ChatGLM Team (2024b), Mistral Jiang et al. (2023). At the same time, we mainly conduct experiments on three categories of benchmarks, and some other capability categories (such as code, reading comprehensive, etc.) also need to be considered to provide more comprehensive and enriching experimental results. In addition, our method still has some flaws in some experiments, and we are consistently studying and optimizing the proposed method.
@@ Appendix A Prompt Templates
Prompt templates in different methods are illustrated in Figure 6, instruct 0-shot and option-ppl are regarded as original methods in our experiments for open-ended tasks and multi-choice tasks respectively.
Figure 6: Prompt templates in different methods

### v2

# BOSE: A Systematic Evaluation Method Optimized for Base Models

@@ Abstract
BOSE: A Systematic Evaluation Method Optimized for Base Models
Hongzhi Luan [^1] Equal contributions.,  Changxin Tianthanks: Equal contributions.,  Zhaoxin Huan  Xiaolu Zhang,  Kunlong Chen,  Zhiqiang Zhang,  Jun Zhouthanks: Corresponding author. Ant Group {luanhongzhi.lhz, tianchangxin.tcx, zhaoxin.hzx}@antgroup.com {yueyin.zxl, kunlong.ckl, lingyao.zzq, jun.zhoujun}@antgroup.com
@@ 1 Introduction
Recently, large language models (LLMs) have demonstrated remarkable achievements across various domains Minaee et al. (2024). This has led to the development of numerous high-performing LLMs OpenAI (2024); Aaron Grattafiori and Abhinav Jauhri (2024); QwenTeam (2025); GemmaTeam (2024). To evaluate the performance of these models on a wide range of tasks, an increasing number of benchmarks have been open-sourced Guo et al. (2023). These benchmarks provide a comprehensive evaluation of the capabilities of LLMs, guiding the training and improving their weakness. Evaluating large language models is a cornerstone in development of LLMs.
@@
To the best of our knowledge, this paper presents the first systematic framework for evaluating base LLMs, introducing empirically validated optimization methodologies with quantitative metrics. We believe this will yield actionable insights and practical recommendations to contribute to advance future base LLM evaluation.
@@ 2 Related Work
Due to the above issues, although many technical reports of open-source LLMs disclose the benchmarks and evaluation results used for base model evaluation Aaron Grattafiori and Abhinav Jauhri (2024); QwenTeam (2025); GemmaTeam (2024), the lack of unified evaluation protocol and the incomplete transparency of evaluation details make it challenging to reproduce and align the evaluation results, which may cause significant difficulties for researchers.
@@ 3.1 Open-ended Task
Light-instruction prompt. While prompting techniques are effective for LLMs Sander Schulhoff and et al. (2024); Wei et al. (2023a), excessive prompting may confuse base models due to their lack of instruction-tuning Wang and Zhou (2024). We propose a lightweight prompt template, which formulates the input using the following format:
@@ 4 Experiments
In this section, we describe systematic experiments to investigate how BOSE enhances evaluation stability during pre-training and ensures consistency between base and instruct models [^1] All of our experiments are implemented based on OpenCompass OpenCompass (2023), and the prompt templates are available in Appendix B..
@@ 4.1.2 Models
Pre-trained models [^3] Since the open-source LLMs do not provide publicly accessible intermediate model weights, we take our pre-trained models for experiments.. These models are collected from our pre-training experiments, encompassing 1B and 2B parameter sizes with different training steps, 96 checkpoints in total, more details about the training recipe can be found in Appendix C and our technical report LingTeam (2025). To assess the capability consistency between base models and their instruction-tuned counterparts, we further fine-tune 24 checkpoints using identical SFT data (size, quality, filtering criteria) and training hyper parameters to derive the instruct models.
Open-source base models. These models mainly cover three remarkable LLM families: (1) Llama family Aaron Grattafiori and Abhinav Jauhri (2024): mainly including Llama-3.1-8B/70B, with an extended analysis incorporating Llama-3-8B/70B and Llama-3.2-1B/3B for consistency investigation; (2) Gemma family GemmaTeam (2024): Gemma-2-9B/27B; (3) Qwen family QwenTeam (2025): Qwen2.5-7B/72B.
@@ 4.1.3 Metric
Consistency with instruct models: measured by the Kendall’s rank correlation between base models’ evaluation scores and corresponding instruct models’ scores, which is inspired by previous work within the community Agarwal et al. (2024).
@@ 4.4 Ablation Study
To investigate whether all components of the proposed ICLiP methodology are indispensable, we conduct ablation experiments here. The following variant prompt templates are considered as comparative methods: (1) instruct_0shot: instruct prompt with 0-shot, also referred to as the original method, (2) instruct_fewshot: instruct prompt with few-shot, (3) light-instruction_0shot: light-instruction prompt with 0-shot, (4) ICLiP, our proposed method. Here, light-instruction refers to the prompt template used in our proposed method, while instruct prompt is the prompt template typically used for instruct model evaluation. Please refer to Appendix A for further implementation details.
@@
To further sourcing the improvements of our proposed method, we conduct two supplementary investigations, and once again proves the validity of the proposed approach. Evaluation details can be found in Appendix B .
@@ 6 Limitations
Our work aims to address some critical issues in base model evaluation, such as instability during pre-training and lack of consistency with instruct models. There may be some subjective biases in our comparative methods, such as the design of instruction prompt templates. As for the considered models, we use our pre-trained checkpoints and 3 open-source LLM families, and there are more open-source LLMs worth exploring, such as DeepSeek DeepSeek-AI (2024), ChatGLM GLMTeam (2024), Mistral Jiang et al. (2023). At the same time, we mainly conduct experiments on three categories of benchmarks, and some other capability categories (such as code, reading comprehensive, etc.) also need to be considered to provide more comprehensive and enriching experimental results. In addition, our method still has some flaws in some experiments, and we are consistently studying and optimizing the proposed method.
Acknowledgments
We sincerely thank the anonymous reviewers for their valuable comments.
@@ Appendix A Prompt Templates
Prompt templates in different methods are illustrated in table 5 to table 10, instruct_0shot and Option-ppl are regarded as original methods in our experiments for open-ended tasks and multi-choice tasks respectively.
PROMPT
Question: {question}
A. {content of optionA}
B. {content of optionB}
C. {content of optionC}
D. {content of optionD}
Answer: A
Table 5: Prompt of Option-PPL, where the underlined part is used for calculating the Perplexity (PPL).
PROMPT
{question} {content of optionA}
Table 6: Prompt of Blank-PPL, where the underlined part is used for calculating the Perplexity (PPL).
PROMPT
You’re a {domain} expert. Given the following question, please reason step by step and put your final answer within boxed{}.
{question}
Table 7: Prompt of Instruct_0shot
PROMPT
You’re a {domain} expert. Given the following question, please reason step by step and put your final answer within boxed{}.
{question_1}
{answer_1}
{question_k}
{answer_k}
{question}
Table 8: Prompt of Instruct_fewshot
PROMPT
Problem: {question}
Solution:
Table 9: Prompt of Light-instruction_0shot
PROMPT
Problem: {question_1}
Solution: {answer_1}
Problem: {question_k}
Solution: {answer_k}
Problem: {question}
Solution:
Table 10: Prompt of ICLiP
Benchmark Instruct_0shot ICLiP
current math_verify upper bound ratio current math_verify upper bound ratio
CMATH 20.58 16.21 22.67 0.715 53.47 52.10 53.91 0.966
Gaokao2023EN 1.04 0.26 4.94 0.053 23.38 23.38 24.42 0.957
CollegeMath 2.91 9.22 9.94 0.928 18.86 23.42 24.41 0.959
Table 11: Ablation study results for sourcing the improvement. $current$: results with current evaluation metrics; $math\_verify$ and $upperbound$: results evaluated with Math_verify and LLM judge respectively; $ratio$: score with math_verify / score with LLM judge, indicating the extent to which the LLM’s predictions follow the format in different methods.
Appendix B Ablation study for sourcing the improvement
We provide some additional ablation studies to further source the improvements of our proposed method. Specifically, with open-ended tasks in consideration, we take Instruct_0shot as a comparative method for simplicity and conduct two supplementary investigations:
i. Comparable evaluation with math_verify [^4] https://github.com/huggingface/Math-Verify. We take the newly released mathematical expression evaluation package math_verify from huggingface to conduct the post-processing and judgment, to identify potential limitations in our current evaluation process.
ii. Loose judgement with LLM as upper bound. To further investigate whether our proposed method effectively ensuring the model’s problem-solving capabilities or simply fitting the answer extraction pattern, we employ GPT-4o to assess the correctness of Base model’s prediction, where only question (NOT prompt) along with prediction and gold answer are provided for judgment and thus eliminate the influence of answer extraction. These scores are regarded as upper-bound evaluation results, with no constraint imposed on the answer format.
As shown in Table 11, we observe that our improvement sources from two aspects:
Better reflection of true capability. While evaluation scores slightly differ across different post-processing methods (relatively bigger difference in the CollegeMath), our proposed method consistently outperforms the comparative methods(i.e., Instruct_0shot).
Higher fitness to the answer extraction. ICLiP exhibits superior alignment with loose scores(upper-bound) compared to both current evaluation process and math-verify judgement, thereby enabling higher tolerance for the post-processing.
The above results further demonstrate the effectiveness of the proposed method.
Appendix C Architecture and Pretraining Setup
We use a GQA (Grouped Query Attention) architecture based on the standard decoder-only Transformer, comprising an embedding layer, alternating layers of attention mechanisms and feed-forward networks. Positional information is handled using RoPE (Rotary Positional Embedding). For training parameters, the pre-trained models are initialized with a standard deviation of $0.006$ and optimized using AdamW, with hyperparameters set to $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=1e^{-8}$ and weight_decay = 0.1. The learning rate follows a WSD (Warm-up, Stabilization, Decay) strategy, where the first $1\%$ of training steps involve linear warm-up. Further architectural and pretraining details are available in Table 12 and our technical report LingTeam (2025).
1B 2B
vocab size 126464 126464
layer num 22 22
hidden size 2048 2560
intermediate size 5632 10240
attention heads 32 40
key value heads 4 8
# Para. (B) 1.487 2.724
# Non-emb. Para. (B) 0.969 2.076
sequence length 4096 4096
learning rate 6.37E-04 5.23E-04
batch size 806 1227
Table 12: The details of architecture and training parameters for pre-trained models in Section 4.1.2